A stock is expected to pay a year-end dividend of $2.00 twelve months from now. The dividend is expected to decline at a rate of 3% a year forever. If the company is in equilibrium and its expected and required rate of return is 17%, which of the following statements is CORRECT?

a.

The company’s current stock price is $14.29.

b.

The company’s expected capital gains yield is 3%.

c.

The company’s expected stock price at the beginning of next year is $9.50.

d.

The constant growth model cannot be used because the growth rate is negative.

e.

The company’s expected stock price at the beginning of next year is $9.70.
Answer #1

The stock price can be found by the Gordon growth formula :

Price = Div/(R - g) =2/(0.17 + 0.03) = $10.

At the beginning of the next year, the stock price will be = 2 x (1-0.03)/(0.17 + 0.03) = $9.7

Hence, only option E is correct.